Course Review for CS 491 A

K. Subramani Department of Computer Science and Electrical Engineering, West Virginia University, Morgantown, WV ksmani@csee.wvu.edu

1 Instructor’s comments

We are halfway through the course and I thought that this is a good time to obtained detailed feedback from you. Your comments are useful for the following reasons:

1. If there is sufficient interest in the subject, I can justify offering it again,

2. If there is too much mathematical content, we can ask the Department to offer courses in Linear Algebra as pre-requisites,

3. I can tailor the second-half of the course to see that your concerns are addressed. It is therefore vital to me that you fill in the rest of this form in as much detail as you can. I did go through some of your responses and noticed that the responses were inadequate. This was partly due to the fact that you were boxed in by the choices! In the current format, you have the freedom to write in exactly what you think of this course and how you would like me to proceed so that you can benefit. I hope you will take this effort seriously.

2 Review

1. Do you like the content of the course?

2. Is the course being covered at the correct pace?

3. Is the course too mathematical?

4. Are you satisfied with the instructor’s help?

5. Do you like the teaching style of the instructor?

6. Would you appreciate similar courses being offered by the Department and/or by this instructor?

7. What are the features about this course that you would like to retain in the second half?

8. What are the feature about this course that you would like to change?

9. What are the attributes of the instructor that you would like to see more of?

10. What are the attributes that you would like changed?

11. Random suggestions